Suppose you throw the 0.12-kg ball at this minimum
Answer #1

You friend is 11 m above you

minimum speed is u. tha ball reaches your friend with zero speed.

using v^2-u^2=2as

0-u^2=-2\times9.81\times11

u=14.69\,m/s

kinetic energy of ball at bottom = Initial kinetic energy of ball + Initial potential energy

\frac{1}{2}mv^2=\frac{1}{2}mu^2+mgh

\frac{1}{2}mv^2=0.12\times9.81\times11+0.12\times9.81\times12=27.1\,J